# MLOmics: Benchmark for Machine Learning on Cancer Multi-Omics Data

### v2

@@ Abstract
Framing the investigation of diverse cancers as a machine learning problem has recently shown significant potential in multi-omics analysis and cancer research. Empowering these successful machine learning models are the high-quality training datasets with sufficient data volume and adequate preprocessing. However, while there exist several public data portals including The Cancer Genome Atlas (TCGA) multi-omics initiative or open-bases such as the LinkedOmics, these databases are not off-the-shelf for existing machine learning models. In this paper we propose MLOmics, an open cancer multi-omics benchmark aiming at serving better the development and evaluation of bioinformatics and machine learning models. MLOmics contains 8,314 patient samples covering all 32 cancer types with four omics types, stratified features, and extensive baselines. Complementary support for downstream analysis and bio-knowledge linking are also included to support interdisciplinary analysis.
@@
Multi-omics analysis has shown great potential to accelerate cancer research. A promising trend consists in framing the investigation of diverse cancers as a machine learning problem, where complex molecular interactions and dysregulations associated with specific tumor cohorts are revealed through integration of multi-omics data into machine learning models. Several impressive achievements have been demonstrated in molecular subtyping [1, 2, 3], disease-gene association prediction [4, 5, 6], and drug discovery [7].
@@
Empowering successful machine learning models are the high-quality training datasets with sufficient data volume and adequate preprocessing. While there exist several public data portals including The Cancer Genome Atlas (TCGA) multi-omics initiative [8] or open-bases such as the LinkedOmics [9], these databases are not off-the-shelf for existing machine learning models. To make these data model-ready, a series of laborious, task-specific processing steps such as metadata review, sample linking, and data cleaning are mandatory. The domain knowledge required, as well as a deep understanding of diverse medical data types and proficiency in bioinformatics tools have become an obstacle for researchers outside of such backgrounds. The gap between the growing body of powerful machine learning models and the absence of well-prepared public data has become a major bottleneck. Currently, some existing research validates proposed machine learning models using inconsistent experimental protocols, including variations in datasets and data processing techniques, and evaluation strategies [10]. These studies could benefit from a fair assessment of extensive baselines on a uniform footing with unified datasets. A benchmark also play a key role in guiding bioinformatics researchers in designing and handpicking the most suitable models.
Figure 1: Schematic workflow of creating the MLOmics. The process starts with collecting patient samples covering 32 cancer types from the TCGA project. All resources in diverse data types and sizes are uniformly integrated and processed to contain data of four omics types. Datasets for benchmark ML tasks were constructed based on the processed data. MLOmics also selected baselines, metrics, and resources to support downstream biological analysis. Overview of the MLOmics. MLOmics provides a user-friendly interface for developing and evaluating machine learning models based on cancer multi-omics data. MLOmics provides datasets in three feature scales for 20 classification, clustering, and omics imputation learning tasks. MLOmics also provides statistical, ML, and DL baselines for each task, which are evaluated by fair metrics. Bio-knowledge database linking with MLOmics. MLOmics provides resources to link with other bio-knowledge databases, enabling integration of outer resources for applications such as ML evaluation, gene-disease association exploration, network inference, and functional analysis.
@@
To meet the growing demand of the community, we introduce MLOmics, an open cancer multi-omics benchmark aiming at serving better the development and evaluation of bioinformatics and machine learning models. MLOmics collected 8,314 patient samples covering all 32 cancer types from the TCGA project. All samples were uniformly processed to contain four omics types: mRNA expression, microRNA expression, DNA methylation, and copy number variations, followed by categorization, protocol verification, feature profiling, transformation, and annotation. Based on the processed data, 20 learning tasks and the associated datasets for pan-cancer analysis, cancer subtypes, and omics imputation were constructed. For each dataset, we provide three feature versions: Original, Aligned, and Top, to support feasible analysis. For example, the Top version contains the most significant features selected via the ANOVA test [11] across all samples to filter out potentially noisy genes. The MLOmics datasets were carefully examined with 6$\sim$10 high citations baseline methods. These baselines were rigorously reproduced and evaluated with various metrics to ensure fair comparison. Complementary resources are included to support basic downstream biological analysis, such as clustering visualization, survival analysis, and Volcano plots. Last but not least, we provide support to interdisciplinary analysis via our locally deployed bio-base resources. Interdisciplinary researchers can retrieve and integrate bio-knowledge of cancer omics studies through resources such as the STRING [12] and KEGG [13]. For instance, exploration of bio-network inference [14] and simulated gene knockouts [15] is supported. In summary, MLOmics provides a user-friendly interface that enables non-experts to easily utilize its open, well-prepared datasets for developing/evaluating machine learning models; conducting interdisciplinary analysis and supporting cancer research and broader biological studies. A detailed overview of the database and its characteristics is provided in Figure. 1.
@@ Methods
Datasets. MLOmics provides three omics datasets with missing values (GS-BRCA, GS-COAD, GS-GBM). Given a full dataset as a matrix $D\in\mathbb{R}^{n\times m}$, we follow prior works [43, 42] to generate a mask matrix $M\in\{0,1\}^{n\times m}$ uniformly at random with the probability of missing $P(M_{ij}=0)=r_{\text{miss}}$, and the probability of retaining $P(M_{ij}=1)=1-r_{\text{miss}}$. The final data matrix is obtained by multiplying element-wise the data matrix $D$ with the mask $M$. The missing level $r_{\text{miss}}$ is selected from $[0.3,0.5,0.7]$.
Baselines. We opt for seven well-recognized methods for imputing missing values, including: Mean Imputation that imputes the missing entry with mean values of entries around it (Mean); K-Nearest Neighbors (KNN) that imputes the missing value with the weighted Euclidean K nearest neighbors; Multivariate Imputation by Chained Equations (MICE) that performs multiple regression to model each missing value conditioned on non-missing values [44]; Iterative SVD (iSVD) that imputes by iterative low-rank SVD decomposition [45]; Spectral Regularization Algorithm (Spectral) that also employs SVD but with a soft threshold and the nuclear norm regularization [46]; Generative Adversarial Imputation Nets (GAIN) that proposes to distinguish between fake and true missing patterns by the generator-discriminator architecture [43]; Graph Neural Network for Tabular Data (GRAPE) that utilizes the graph networks to impute based on learned information from columns and rows of the data matrix [42].
Metrics. We use the Mean Squared Error (MSE) between the unmasked entries ($M_{ij}=1$) as the training loss to let the model predict the actual value. During test, the masked missing values are used for evaluating the model performance ($M_{ij}=0$).
MLOmics will be continuously updated with baselines and evaluations on the defined learning tasks. Detailed description of the baselines and metrics is provided in the Supplementary Material.
Biological Database Linking and Downstream Analysis
A rising trend in multi-omics analysis is to integrate multi-omics data (non-network data) with biological networks to better understand complex functions on the gene or protein level. MLOmics provides offline linking resources for well-established databases such as STRING [12] and KEGG [13]. STRING offers network structural information about genes, while KEGG provides systematic functional information on various biological networks.
To integrate omics data into biological networks, omics features (e.g., genes) are first mapped to STRING and KEGG network nodes (e.g., gene products) using standardized identifiers, e.g. unified gene IDs. MLOmics provides mapping files to extract shared gene IDs and standardize inconsistent file formats, creating a unified feature list. This feature list includes shared gene IDs along with corresponding omics features or measurements. Once the omics features are successfully mapped, they can be incorporated into networks by assigning weights to nodes or edges based on the data e.g. expression levels. Additionally, MLOmics provides a suite of biological analysis tools to evaluate the significance of results generated by different machine learning models. These tools include survival analysis [47], gene differential expression (GDE) analysis [48], featuring volcano plots and simulated knockout analyses at the single-gene level, and KEGG pathway analysis [49], which operates at the gene function set level. These resources are designed to support researchers in conducting classic and reliable biological validations of their data analysis frameworks.
@@
As shown in Fig. 2, MLOmics framework consists of three major components: (1) main datasets that includes all cancer multi-omics datasets for various tasks; (2) baselines and metrics that provides the source code for baseline models and performance metrics and (3) downstream analysis tools and reources linking that provides tools for further analysis and links to additional biological resources.
Figure 2: Schematic of MLOmics resources structure. MLOmics framework consists of three major components. The Main Datasets file includes all cancer multi-omics datasets for various tasks. The Baseline and Metrics file provides the source code for baseline models and performance metrics. The Downstream Analysis Tools and Resources Linking file offers sources for further analysis and links to additional biological resources.
Figure 3: The experiment results and downstream analysis of applying machine learning baselines based on MLOmics datasets. (a) The PREC bar plot for each baseline method across all datasets. (b) The SIL heatmap for each baseline method across all datasets. (c) The box plot for each baseline method on three imputation datasets. (d-e) The downstream analysis results based on the clustering outcomes of XOmiVAE applied to the BRCA patient clustering datasets.
Main Datasets
The Main Datasets repository is stored as csv files and is organized into three layers. The first layer contains three files corresponding to three different tasks: Classification_datasets, Clustering_datasets, and Imputation_datasets. The second layer includes files for specific tasks, such as GS-BRCA, ACC, and Imp-BRCA. The third layer contains three files corresponding to different feature scales, i.e., Original, Aligned, and Top. The omics data from different omics sources are stored in the following files: Cancer_miRNA_Feature.csv, Cancer_mRNA_Feature.csv, Cancer_CNV_Feature.csv, and Cancer_Methy_Feature.csv. Here, Cancer represents the cancer type, and Feature indicates the feature scale type. The ground truth labels are provided in the file Cancer_label_num.csv, where Cancer represents the cancer type. The patient survival records are stored in the file Cancer_survival.csv.
Baselines and Metrics
The Baselines and Metrics repository contains .py (Python) and .r files and is organized into following three layers: the first layer contains three folders corresponding to three study categories: Classification, Clustering, and Imputation. The second layer contains two folders Metrics and Baselines corresponding to the metrics code and baseline code. The third layer of Baselines contains .py and .r files that are the actual implementation of baselines. The metrics code is provided in the file Study_metrics.py, where Study represents the three study types. The baseline code is provided in the file Baseline.py or Baseline.r, where Baseline represents the name of a specific baseline.
Downstream Analysis Tools and Resources Linking
The Downstream Analysis Tools and Resources Linking repository contains .csv, .py, .r files. It comprises three folders Knowledge_bases, Clinical_annotation, and Analysis_tools. In the knowledge base repository Knowledge_bases, the mapping files to databases are provided in the file DB.csv, where DB represents the name of a specific database, such as KEGG. In Clinical_annotation, the clinical annotations of patients are provided in the file Clinical_Rec.csv. In Analysis_tools, the downstream analysis code is provided in the file Analysis.py or Analysis.r, where Analysis represents the name of a specific analysis tool, such as Pathway.
@@
To assess the potential utility of the MLOmics datasets for machine learning studies on cancer multi-omics, we first evaluated baseline methods on the MLOmics datasets for classification, clustering, and omics data imputation tasks. Next, we performed downstream analysis to interpret the biological significance of the results. The experiment results are summarized in Figure. 3.
Figure. 3(a) shows the PREC bar plot for each baseline method across all datasets. Overall, machine learning-based methods (Subtype-GAN, DCAP, MAUI, XOmiVAE, CustOmics, and DeepCC) outperformed traditional statistical methods (SVM, XGBoost, RF, and LR). The performance of traditional methods was relatively uniform, whereas notable gaps were observed among the machine learning-based methods. These results highlight the significant potential of machine learning-based approaches in cancer patient classification tasks.
Figure 3(b) shows the SIL heatmap for each baseline method across all datasets. Overall, methods with deep generative neural network architectures (Subtype-GAN, DCAP, MAUI, XOmiVAE, and MCluster-VAEs) outperformed the others (SNF, NEMO, CIMLR, iClusterBayes, and moCluster). Compared to the results from patient classification tasks, baseline methods for clustering tasks were more sensitive to dataset characteristics (such as sample size). This finding suggests that cancer subtyping is more complex than cancer classification, which aligns with the consensus in the cancer research community.
Figure 3(c) shows the box plot for each baseline method on three imputation datasets (i.e., Imp-BRCA, Imp-COAD, and Imp-GBM). Overall, matrix decomposition methods (SVD, Spectral) have performed better than deep learning-based methods (GAIN, GNN). This suggests that these methods might have captured inherent properties and the low-rank nature of the data, demonstrating the potential of matrix decomposition methods in imputation. This result indicates that traditional matrix decomposition methods still perform well in predicting missing omics values, whereas deep learning-based methods have room for improvement.
Figure 3(d) and (e) present the downstream analysis results based on the clustering outcomes of XOmiVAE applied to the BRCA patient clustering datasets. In Figure 3(d), the first subfigure displays the Kaplan-Meier survival curves, illustrating the survival differences across patient clusters. The second subfigure presents a volcano plot of differentially expressed genes between two patient clusters, highlighting representative genes such as MAP2K2 and PIK3CA. The third subfigure shows a UMAP visualization of patient samples in the latent space, with colors clearly distinguishing different patient groups. In Figure 3(e), the first subfigure displays a partial example of a KEGG pathway plot, where upregulated and downregulated genes between two clusters are shown in red and blue, respectively. These genes are mapped onto the HER2 and Basal-like pathways, highlighting key regulatory locations linking to the KEGG database. The second subfigure presents gene networks of the differentially expressed genes, based on mapping files linked to the STRING database. It reveals distinct structural differences between the two gene networks. The third subfigure visualizes patient samples before and after a simulated gene knockout, demonstrating a notable shift. This shift suggests that the differentially expressed genes may contribute to transitions between the two patient groups. Overall, these downstream analyses demonstrate the biological relevance and feasibility of the baseline results, including evaluating outcomes through survival analysis and differential gene expression analysis and conducting further explorations with the aid of advanced biological databases and approaches.
@@
Figure 4 illustrates the usage of MLOmics, including setup instructions, code usage for custom model-task-dataset specifications, and downstream analysis workflows.
Using specific datasets for ML tasks
After successfully setting up MLOmics, use the following command to select a specific dataset to train model on:
./<baseline_model>.sh <dataset> <version> [options]
Where:
  • <baseline_model>: Name of the model script (e.g., GRAPE.sh).
  • <dataset>: Target dataset name (e.g., GS-BRCA).
  • <version>: Feature version (e.g., Original, Aligned, Top).
  • [options]: Optional parameters such as missing rate (e.g., 0.3).
Below are examples of MLOmics parameter settings for classification, clustering, and imputation tasks:
  • Classification:
    # Select the original scale GS-BRCA dataset
    # and train the DeepCC model for the classification task.
    cd Scripts/Classification
    ./DeepCC.sh GS-BRCA Original
  • Clustering:
    # Select the top scale ACC dataset
    # and train the Subtype-GAN model for the clustering task.
    cd Scripts/Clustering
    ./Subtype-GAN.sh ACC Top
  • Imputation:
    # Select the top scale ACC dataset
    # and train the GAIN model for the imputation task (0.3 missing rate).
    cd Scripts/Imputation
    ./GAIN.sh ACC Top 0.3
Downstream Analysis
After obtaining classification or clustering results, use the following command to process downstream gene expression and pathway enrichment analyses:
  • Volcano Plot Generation:
    # Differential gene expression analysis (P-value = 0.05) with volcano plot.
    cd Scripts/Dwonstream_Analysis
    ./volcano.sh <clustering_log_path> [options]
    --p_value_threshold 0.05
  • KEGG Pathway Analysis:
    # Pathway enrichment analysis (P-value = 0.05) with
    # KEGG Pathway visualization.
    cd Scripts/Dwonstream_Analysis
    ./pwanalysis.sh <clustering_log_path> [options]
    --p_value_cutoff 0.05
Where:
  • <clustering_log_path>: Location of resultant sample label.
  • [options]: Optional parameters such as P-value (e.g., p_value_threshold 0.05).
@@
The MLOmics is publicly available on GitHub under the Creative Commons 4.0 Attribution (CC-BY-4.0). All code and resources are stored on the cloud in user-friendly formats and are accessible via our GitHub repository (https://github.com/chenzRG/Cancer-Multi-Omics-Benchmark). We provide comprehensive guidelines for utilization. All files are ready for direct loading and analysis using standard Python data packages like Numpy or Pandas. We hope it can lower the barriers to entry for machine learning researchers interested in developing methods for cancer multi-omics data analysis, thereby encouraging rapid progress in the field.

### v3

# MLOmics: Cancer Multi-Omics Database for Machine Learning

@@ Abstract
Framing the investigation of diverse cancers as a machine learning problem has recently shown significant potential in multi-omics analysis and cancer research. Empowering these successful machine learning models are the high-quality training datasets with sufficient data volume and adequate preprocessing. However, while there exist several public data portals, including The Cancer Genome Atlas (TCGA) multi-omics initiative or open-bases such as the LinkedOmics, these databases are not off-the-shelf for existing machine learning models. In this paper, we introduce MLOmics, an open cancer multi-omics database aiming at serving better the development and evaluation of bioinformatics and machine learning models. MLOmics contains 8,314 patient samples covering all 32 cancer types with four omics types, stratified features, and extensive baselines. Complementary support for downstream analysis and bio-knowledge linking are also included to support interdisciplinary analysis.
@@
Multi-omics analysis has shown great potential to accelerate cancer research. A promising trend consists of framing the investigation of diverse cancers as a machine learning problem, where complex molecular interactions and dysregulations associated with specific tumor cohorts are revealed through integration of multi-omics data into machine learning models. Several impressive achievements have been demonstrated in molecular subtyping [1, 2, 3], disease-gene association prediction [4, 5, 6], and drug discovery [7].
@@
Empowering successful machine learning models are the high-quality training datasets with sufficient data volume and adequate preprocessing. While there exist several public data portals including The Cancer Genome Atlas (TCGA) multi-omics initiative [8] or open-bases such as the LinkedOmics [9], these databases are not off-the-shelf for existing machine learning models. To make these data model-ready, a series of laborious, task-specific processing steps such as metadata review, sample linking, and data cleaning are mandatory. The domain knowledge required, as well as a deep understanding of diverse medical data types and proficiency in bioinformatics tools have become an obstacle for researchers outside of such backgrounds. The gap between the growing body of powerful machine learning models and the absence of well-prepared public data has become a major bottleneck. Currently, some existing methods validate their proposed machine learning models using inconsistent experimental protocols, with variations in datasets, data processing techniques and evaluation strategies [10]. These studies could benefit from a fair assessment of extensive baselines on a uniform footing with unified datasets and to offer more reliable recommendations on models.
@@
To meet the growing demand of the community, we introduce MLOmics, an open cancer multi-omics database aiming at serving better the development and evaluation of bioinformatics and machine learning models. MLOmics collected 8,314 patient samples covering all 32 cancer types from the TCGA project. All samples were uniformly processed to contain four omics types: mRNA expression, microRNA expression, DNA methylation, and copy number variations, followed by categorization, protocol verification, feature profiling, transformation, and annotation. For each dataset, we provide three feature versions: Original, Aligned, and Top, to support feasible analysis. For example, the Top version contains the most significant features selected via the ANOVA test [11] across all samples to filter out potentially noisy genes. The MLOmics datasets were carefully examined with 6$\sim$10 highly cited baseline methods. These baselines were rigorously reproduced and evaluated with various metrics to ensure fair comparison. Complementary resources are included to support basic downstream biological analysis, such as clustering visualization, survival analysis, and Volcano plots. Last but not least, we provide support to interdisciplinary analysis via our locally deployed bio-base resources. Interdisciplinary researchers can retrieve and integrate bio-knowledge of cancer omics studies through resources such as the STRING [12] and KEGG [13]. For instance, exploration of bio-network inference [14] and simulated gene knockouts [15] is supported. In summary, MLOmics is an open, unified database approachable to non-experts for developing/evaluating machine learning models; conducting interdisciplinary analysis and supporting cancer research and broader biological studies. We showcase the above usages in the Usage Notes section. A detailed overview of the MLOmics database and its characteristics is provided in Figure 1.
@@ Methods
Datasets: MLOmics provides five omics datasets with missing values (Imp-BRCA, Imp-COAD, Imp-GBM, Imp-LGG, and Imp-OV). Given a full dataset as a matrix $D\in\mathbb{R}^{n\times m}$, we follow prior works [43, 42] to generate a mask matrix $M\in\{0,1\}^{n\times m}$ uniformly at random with the probability of missing $P(M_{ij}=0)=r_{\text{miss}}$, and the probability of retaining $P(M_{ij}=1)=1-r_{\text{miss}}$. The final data matrix is obtained by multiplying element-wise the data matrix $D$ with the mask $M$. The missing level $r_{\text{miss}}$ is selected from $[0.3,0.5,0.7]$.
Baselines: We opt for seven well-recognized methods for imputing missing values, including: Mean Imputation that imputes the missing entry with mean values of entries around it (Mean); K-Nearest Neighbors (KNN) that imputes the missing value with the weighted Euclidean K nearest neighbors; Multivariate Imputation by Chained Equations (MICE) that performs multiple regression to model each missing value conditioned on non-missing values [44]; Iterative SVD (iSVD) that imputes by iterative low-rank SVD decomposition [45]; Spectral Regularization Algorithm (Spectral) that also employs SVD but with a soft threshold and the nuclear norm regularization [46]; Generative Adversarial Imputation Nets (GAIN) that proposes to distinguish between fake and true missing patterns by the generator-discriminator architecture [43]; Graph Neural Network for Tabular Data (GRAPE) that utilizes the graph networks to impute based on learned information from columns and rows of the data matrix [42].
Metrics: We use the Mean Squared Error (MSE) between the unmasked entries ($M_{ij}=1$) as the training loss to let the model predict the actual value. During the test, the masked missing values are used for evaluating the model performance ($M_{ij}=0$).
MLOmics will be continuously updated with baselines and evaluations on the defined learning tasks. A detailed description of the baselines and metrics is provided in the Supplementary Material.
@@
The MLOmics main datasets [47] are available on Figshare (https://figshare.com/articles/dataset/MLOmics_Cancer_Multi-Omics_Database_for_Machine_Learning/28729127) and Hugging Face (https://huggingface.co/datasets/AIBIC/MLOmics). The MLOmic main datasets are now accessible under the Creative Commons 4.0 Attribution (CC-BY-4.0), which supports its use for educational and research purposes. The main datasets presented include all cancer multi-omics datasets corresponding to the various tasks described above.
MLOmics Overview
The MLOmics repository is structured into three sections: (1) Main Datasets, (2) Baseline and Metrics, and (3) Downstream Analysis Tools and Resources Linking.
The core section is the (1) Main Datasets: the repository hosts a comprehensive collection of tasks-ready cancer multi-omics datasets, stored primarily as .csv (comma-separated values) files.
Along with the other two sections:(2) Baseline and Metrics: the repository provides source codes of baseline models and evaluation metrics for different tasks, typically implemented in Python or R code; (3) Downstream Analysis Tools and Resources Linking: the repository encompasses additional tools and resources that complement the main datasets and downstream omics analysis needs, implemented as .csv, .py or R files.
In the following parts of this section, we focus on presenting the properties of the Main Datasets, detailing its organizational structure, principal components, and resources.
Main Datasets Format
Main datasets of MLOmics use .csv files to manage and store all omics datasets, widely favored in biomedical research, including multi-omics studies. Despite its simplicity, .csv remains efficient even with large datasets encountered in genomic and proteomic studies. Both Python and R provide built-in functions and libraries to read, write, and analyze .csv files efficiently.
Specifically, MLOmics provides its main datasets in .csv files as plain-text files, where commas separate data values. These files maintain a structured and intuitive format, with rows representing individual data records and columns corresponding to different attributes or variables. The key characteristics of the main dataset files are:
  • Encoding: UTF-8, ensuring broad compatibility across different systems and software.
  • Delimiter: , (comma), used to separate values in each row.
[…]
  • Header Row: The first row contains column names, typically representing sample identifiers (e.g., Sample1, Sample2, Sample3, etc.).
  • Data Rows: Each subsequent row corresponds to a distinct feature, such as a gene, with attributes like expression values recorded across multiple patient samples.
  • Values: Numeric values (float), representing continuous measurements of omics attributes.
For example, consider a simplified .csv file containing mRNA data for a set of gene features (rows) across several patient samples (columns):
Feature Sample1 Sample2 Sample3 Sample4
GeneA 0.23 0.18 0.35 0.21
GeneB 0.56 0.49 0.52 0.58
GeneC 0.19 0.22 0.15 0.17
GeneD 0.08 0.10 0.09 0.12
In this example, each row corresponds to a specific gene (GeneA, GeneB, GeneC, GeneD). Each column represents a different sample (Sample1, Sample2, Sample3, Sample4). The numeric values (variables) in the cells denote the expression levels of each gene in each sample.
Main Datasets Structure
As shown in Fig. 2, the main datasets repository is organized into three layers:
(1) The first layer contains three files corresponding to three different tasks: Classification_datasets, Clustering_datasets, and Imputation_datasets.
(2) The second layer includes files for specific tasks, such as GS-BRCA, ACC, and Imp-BRCA.
(3) The third layer contains three files corresponding to different feature scales, i.e., Original, Aligned, and Top.
The omics data from different omics sources are stored in the following files: Cancer_miRNA_Feature.csv, Cancer_mRNA_Feature.csv, Cancer_CNV_Feature.csv, and Cancer_Methy_Feature.csv. Here, Cancer represents the cancer type, and Feature indicates the feature scale type.
The ground truth labels are provided in the file Cancer_label_num.csv, where Cancer represents the cancer type. The patient survival records are stored in the file Cancer_survival.csv.
@@
The Methods section detailed the data collection, preprocessing and curation of the MLOmics datasets. To validate the collected datasets, we set up a series of classification, clustering and imputation experiments each with a wide array of models ranging from conventional statistical models to deep neural networks. The experimental results and evaluations are summarized in Figure 3.
@@
We provide comprehensive guidelines for utilizing the MLOmics repository, as illustrated in Figure 4. All codes and files are ready for direct loading and analysis using standard Python data packages such as NumPy and Pandas. As a demonstration of how to use MLOmics for machine learning applications, we provide 20 classification, clustering, and imputation tasks with fair evaluation protocols for pan-cancer analysis, cancer subtypes, and omics imputation. Access to these resources is provided in the Code Availability section. A rising trend in multi-omics analysis is to integrate multi-omics data (non-network data) with biological networks to better understand complex functions on the gene or protein level. MLOmics provides offline linking resources for well-established databases such as STRING [12] and KEGG [13]. We hope it can lower the barriers to entry for machine learning researchers interested in developing methods for cancer multi-omics data analysis, thereby encouraging rapid progress in the field.
@@
All codes and resources of MLOmics are publicly available under the CC-BY-4.0. The code for preprocessing and preparing the main MLOmics databases, as well as the benchmarking algorithms used in MLOmics, is available at the following repository: (https://github.com/chenzRG/Cancer-Multi-Omics-Benchmark).